# Fisher Flow Matching for Generative Modeling over Discrete Data

### v3

@@ Abstract
Generative modeling over discrete data has recently seen numerous success stories, with applications spanning language modeling, biological sequence design, and graph-structured molecular data. The predominant generative modeling paradigm for discrete data is still autoregressive, with more recent alternatives based on diffusion or flow-matching falling short of their impressive performance in continuous data settings, such as image or video generation. In this work, we introduce Fisher-Flow, a novel flow-matching model for discrete data. Fisher-Flow takes a manifestly geometric perspective by considering categorical distributions over discrete data as points residing on a statistical manifold equipped with its natural Riemannian metric: the Fisher-Rao metric. As a result, we demonstrate discrete data itself can be continuously reparameterised to points on the positive orthant of the $d$-hypersphere $\mathbb{S}^{d}_{+}$, which allows us to define flows that map any source distribution to target in a principled manner by transporting mass along (closed-form) geodesics of $\mathbb{S}^{d}_{+}$. Furthermore, the learned flows in Fisher-Flow can be further bootstrapped by leveraging Riemannian optimal transport leading to improved training dynamics. We prove that the gradient flow induced by Fisher-Flow is optimal in reducing the forward KL divergence. We evaluate Fisher-Flow on an array of synthetic and diverse real-world benchmarks, including designing DNA Promoter, and DNA Enhancer sequences. Empirically, we find that Fisher-Flow improves over prior diffusion and flow-matching models on these benchmarks.
@@ 2.2 Flow matching over Riemannian manifolds
A probability path on a Riemannian manifold, ${\mathcal{M}}^{d}$, is a continuous interpolation between two distributions, $p_{0},p_{1}\in{\mathcal{P}}({\mathcal{M}}^{d})$, indexed by time $t$. Let $p_{t}$ be a distribution on a probability path that connects $p_{0}$ to $p_{1}$ and consider its associated flow, $\psi_{t}$, and vector field, $u_{t}$. We can learn a continuous normalising flow (CNF) by directly regressing the vector field, $u_{t}$, with a parametric one, $v_{\theta}\in{\mathcal{T}}{\mathcal{M}}^{d}$, where ${\mathcal{T}}{\mathcal{M}}^{d}$ is the tangent bundle. In effect, the goal of learning is to match the flow—termed flow-matching—of the target vector field and can be formulated into a simulation-free training objective [45, FM], provided $p_{t}$ satisfies the boundary conditions, $p_{0}=p_{\rm{data}}$ and $p_{1}=p_{\text{prior}}$. As stated, the vanilla flow matching objective is intractable as we generally do not have access to the closed-form of $u_{t}$ that generates $p_{t}$. Instead, we can opt to regress $v_{\theta}$ against a conditional vector field, $u_{t}(x_{t}|z)$, generating a conditional probability path $p_{t}(x_{t}|z)$, and use it to recover the target unconditional path: $p_{t}(x_{t})=\int_{{\mathcal{M}}}p_{t}(x_{t}|z)q(z)\mathrm{d}z$. Similarly, the vector field $u_{t}$ can also be recovered by marginalising conditional vector fields, $u_{t}(x|z)$. This allows us to state the CFM objective for Riemannian manifolds [25]:
@@ 4.1 Synthetic experiments
Density learning in arbitrary dimensions. We also consider the toy experiment of Stark et al. [60], where we seek to model a random distribution over $\Delta_{1}^{K}\times\cdots\times\Delta_{4}^{K}$ for varying $K\in\mathbb{N}^{\star}$, To evaluate the model, we report the KL divergence of the estimated distribution over $512{,}000$ samples, $\tilde{p}$, and the true generated distribution, $q$, i.e., $\mathbb{D}_{\rm KL}(\tilde{p}\mid\mid q)$. Further details are provided in Section F.4.1. Results in Figure 3(b) demonstrate that Fisher-Flow outperforms Dirichlet FM, which instead uses a Dirichlet conditional path on $\Delta^{d}$. We also conduct an ablation in Figure 3(a) and find that using optimal transport helps for both Fisher-Flow on $\mathring{\Delta}^{d}$ and $\mathbb{S}_{+}^{d}$, with the latter leading to the best performance.
@@
(a) Results of the ablation study.
(b) KL divergence of Fisher-Flow against DFM.
@@ 4.2 Promoter DNA sequence design
Table 1: MSE of the transcription profile conditioned on generated promoter DNA sequences over the test set. We obtained the final $3$ MSE and perplexity (PPL) values by running $5$ independent experiments. All other baseline numbers are taken directly from Stark et al. [60].
@@
Results. Our experimental evaluation closely follows prior work [9, 60]. We report the MSE between the signal of our conditionally generated sequence and the target one, a human genome promoter sequence (MSE in Table 1), both given by the same pre-trained Sei model [24]. We train our model on $88{,}470$ promoter sequences, each of length $1{,}024$, from a database of human promoters [39], each sequence having an associated expression level indicating the likelihood of transcription at each DNA position. As shown in Table 1, Fisher-Flow outperforms baseline methods DDSM [9] and DFM [60] on the MSE evaluation. Perplexities (PPL) from Fisher-Flow on the test set are also better than the baselines such as an autoregressive model and, on average, improve over DFM.
4.3 Enhancer DNA design
@@
Enhancers are DNA sequences that regulate the transcription of DNA in specific cell types, e.g., melanoma cells. Prior work has made use of generative models for designing enhancer DNA sequences in specific cells [63]. Following Stark et al. [60], we report the perplexity over sequences as the main measure of performance. We also include results on Fréchet Biological Distance (FBD) with pre-trained classifiers provided in DFM [60], cf.  §F.4.3. Nevertheless, those classifiers perform poorly on cell-type classification of Enhancer sequences, with test set accuracies of $11.5\%$ and $11.2\%$ on the Melanoma and FlyBrain datasets, respectively; thus, metrics derived from these are not representative of model quality, which we still included in §F.4.3 for transparency.
@@
Results. We report our results in Table 2, with FBD reported in Table 3. We observe that Fisher-Flow obtains significantly better performance than DFM, which highlights its ability to fit the distribution of Melanoma and FlyBrain DNA enhancer sequences. Moreover, we also note that our method improves over the language model baseline on both datasets, which bolsters the belief that Fisher-Flow can be used in similar settings to those of autoregressive models.
@@ Acknowledgements
OD is supported by both Project CETI and Intel. MP is supported by CenIA and by Chilean Fondecyt grant n. 1210426. AJB is partially supported by an NSERC Post-doc fellowship. This work was also partially funded by EPSRC Turing AI World-Leading Research Fellowship No. EP/X040062/1.
@@ Appendix B Geometry of the Simplex
    Then a restatement of Chentsov’s theorem [66, Thm. 11.1], [11, Thm. 1.2] is that if a sequence of Riemannian metrics $g_{d}$ over $\Delta^{d}$ defined for $d\geq 2$ satisfies the property that for any $1\leq d^{\prime}\leq d$ any Markov morphism $M:\Delta^{d^{\prime}}\to\Delta^{d}$ is an isometry with respect to metrics $g_{d^{\prime}},g_{d}$, then there exists $C>0$ such that each of the $g_{d}$ is $C$ times the Fisher-Rao metric on $\Delta^{d}$.
@@
For further details see Villani et al. [68, Ch. 13].
@@
Step 1. Note that by a small modification of the proof, we can apply Villani et al. [68, Thm. 10.42] to $\Delta^{d}$ with cost equal to $\mathbb{D}_{\rm{KL}}$, and obtain that the $W_{\rm{KL}}$-distance between an admissible competitor $p(\theta+\delta\theta)$ in Eq. 15 and $p(\theta_{n})$ is realised by a transport plan $T^{\delta\theta}$, such that we have $p(\theta+\delta\theta)=T^{\delta\theta}_{\#}p(\theta)$. By definition of $W_{\rm{KL}}$ and due to Chebyshev’s inequality, for all $C>0$, the set of points $S_{C}$ that $T^{\delta\theta}$ moves by more than $C\epsilon$ in $\mathbb{D}_{\rm{KL}}$-distance has $p(\theta)$-measure not larger than $1/C$. Furthermore, $T^{\delta\theta}$ is uniformly bounded over $\Delta^{d}_{c}\setminus S_{C}$ by our initial hypothesis. By approximating this transport plan by a flow (one can adapt the ideas from e.g., Santambrogio [57, Thm. 4.4] for this contstruction) over $S_{C}$, we can find a vector field $v^{\delta\theta}$ such that $v^{\delta\theta}(p_{\omega})=\frac{1}{\epsilon}\log_{p_{\omega}}(T^{\delta\theta}(p_{\omega}))+o_{\epsilon}(|\delta\theta|)$ for $p_{\omega}\in S_{C}$, with error uniformly bounded in $p_{\omega}\in{\mathcal{M}}$. We then extend $v^{\delta\theta}$ arbitrarily outside $S_{C}$. This procedure associates to each small enough change $\delta\theta$ a vector field $v_{\delta\theta}\in T_{p(\theta)}{\mathcal{P}}$ which whose time-$\epsilon$ flow, denoted $\phi_{v_{\delta\theta}}(t=\epsilon,\cdot)$ pushes measure $p(\theta)$ to a measure approaching $p(\theta+\delta\theta)$ in the limit $\epsilon\to 0,C\to\infty$.
@@ Appendix D Optimal Transport proofs
For point 1, we can use Villani et al. [68, Thm. 10.28] (the simpler Villani et al. [68, Thm. 10.41] also applies, with the minor modification that we work on a manifold with boundary). To verify its conditions, note that ${\mathcal{M}}^{d}\subset\mathbb{S}_{+}$ is a subset of a Riemannian manifold and has $(d-1)$-dimensional measure, and that cost $c(x,y)=d^{2}(x,y)$ is convex, thus it has unique superdifferential and $\nabla_{x}c(x,\cdot)$ is injective, as required.
For points 2 and 3, we note that by Villani et al. [68, Cor. 7.22] (see also McCann [51]), in general Polish spaces displacement interpolants as given by Eq. 19 and Eq. 20, coincide with Wasserstein geodesics.
A simplified version of the proof of 4. is present in Santambrogio [57, Prop. 5.30]. For the general case, we can use Villani et al. [68, Thm. 10.28], in particular eq. (10.20) therein. Note that for $c(x,y)=d^{2}(x,y)$, as indicated in Example 10.36 this equation corresponds to the equation of geodesics in the underlying manifold. Then we just note that $u_{t}$ is the velocity field of a constant speed geodesic.
Point 5 is a special case of Villani et al. [68, Thm. 7.21], see also Granieri [32].
@@ E.1 Dirichlet Flow matching
where $\tilde{I}_{x}(a,b)=\frac{\partial}{\partial a}I_{x}(a,b)$ is the derivative of the regularised incomplete beta function [60, Appendix A.1] and $C\propto 1/t$ as in regular linear flow matching.

### v4

@@ Abstract
Generative modeling over discrete data has recently seen numerous success stories, with applications spanning language modeling, biological sequence design, and graph-structured molecular data. The predominant generative modeling paradigm for discrete data is still autoregressive, with more recent alternatives based on diffusion or flow-matching falling short of their impressive performance in continuous data settings, such as image or video generation. In this work, we introduce Fisher-Flow, a novel flow-matching model for discrete data. Fisher-Flow takes a manifestly geometric perspective by considering categorical distributions over discrete data as points residing on a statistical manifold equipped with its natural Riemannian metric: the Fisher-Rao metric. As a result, we demonstrate discrete data itself can be continuously reparameterised to points on the positive orthant of the $d$-hypersphere $\mathbb{S}^{d}_{+}$, which allows us to define flows that map any source distribution to target in a principled manner by transporting mass along (closed-form) geodesics of $\mathbb{S}^{d}_{+}$. Furthermore, the learned flows in Fisher-Flow can be further bootstrapped by leveraging Riemannian optimal transport leading to improved training dynamics. We prove that the gradient flow induced by Fisher-Flow is optimal in reducing the forward KL divergence. We evaluate Fisher-Flow on an array of synthetic and diverse real-world benchmarks, including designing DNA Promoter, and DNA Enhancer sequences. Empirically, we find that Fisher-Flow improves over prior diffusion and flow-matching models on these benchmarks. Our code is available at https://github.com/olsdavis/fisher-flow.
@@ 2.2 Flow matching over Riemannian manifolds
A probability path on a Riemannian manifold, ${\mathcal{M}}^{d}$, is a continuous interpolation between two distributions, $p_{0},p_{1}\in{\mathcal{P}}({\mathcal{M}}^{d})$, indexed by time $t$. Let $p_{t}$ be a distribution on a probability path that connects $p_{0}$ to $p_{1}$ and consider its associated flow, $\psi_{t}$, and vector field, $u_{t}$. We can learn a continuous normalising flow (CNF) by directly regressing the vector field, $u_{t}$, with a parametric one, $v_{\theta}\in{\mathcal{T}}{\mathcal{M}}^{d}$, where ${\mathcal{T}}{\mathcal{M}}^{d}$ is the tangent bundle. In effect, the goal of learning is to match the flow—termed flow-matching—of the target vector field and can be formulated into a simulation-free training objective [48, FM], provided $p_{t}$ satisfies the boundary conditions, $p_{0}=p_{\rm{data}}$ and $p_{1}=p_{\text{prior}}$. As stated, the vanilla flow matching objective is intractable as we generally do not have access to the closed-form of $u_{t}$ that generates $p_{t}$. Instead, we can opt to regress $v_{\theta}$ against a conditional vector field, $u_{t}(x_{t}|z)$, generating a conditional probability path $p_{t}(x_{t}|z)$, and use it to recover the target unconditional path: $p_{t}(x_{t})=\int_{{\mathcal{M}}}p_{t}(x_{t}|z)q(z)\mathrm{d}z$. Similarly, the vector field $u_{t}$ can also be recovered by marginalising conditional vector fields, $u_{t}(x|z)$. This allows us to state the CFM objective for Riemannian manifolds [26]:
@@ 4.1 Synthetic experiments
Density learning in arbitrary dimensions. We also consider the toy experiment of Stark et al. [68], where we seek to model a random distribution over $(\Delta^{K})^{4}$ for $K\in\mathbb{N}^{\star}$. TWhe KL divergence between the estimated distribution over $512{,}000$ samples and the true generated distribution is used as the evaluation metric. Details are provided in §F.4.1. Results in Figure 3(b) demonstrate that Fisher-Flow outperforms Dirichlet FM, while remaining competitive against D3PM [8] and Multinomial Flow [41], especially in high dimensions, in which both exhibit unstable behaviour. We also conduct an ablation in Figure 3(a) and find that using optimal transport helps for both Fisher-Flow on $\mathring{\Delta}^{d}$ and $\mathbb{S}_{+}^{d}$, with the latter leading to the best performance.
@@
(a) Results of the ablation study. Missing points for Multinomial Flow [41] and D3PM [8] are NaNs.
(b) KL divergence of Fisher-Flow against Dirichlet FM.
@@ 4.2 Promoter DNA sequence design
Table 1: MSE of the transcription profile conditioned on generated promoter DNA sequences over the test set. The last $3$ MSE and PPL values are from $5$ independent experiments. The remaining numbers are taken directly from Stark et al. [68].
@@
Results. Our experimental evaluation closely follows prior work [9, 68]. We report the MSE between the signal of our conditionally generated sequence and the target one, a human genome promoter sequence (MSE in Table 2), both given by the same pre-trained Sei model [25]. We train our model on $88{,}470$ promoter sequences, each of length $1{,}024$, from a database of human promoters [40], each sequence having an associated expression level indicating the likelihood of transcription at each DNA position. As shown in Table 2, Fisher-Flow outperforms baseline methods DDSM [9] and Dirichlet FM [68] on the MSE evaluation. Perplexities (PPL) from Fisher-Flow on the test set are also better than the baselines and, on average, improve over Dirichlet FM.
4.3 Enhancer DNA design
@@
Enhancers are DNA sequences that regulate the transcription of DNA in specific cell types (e.g., melanoma cells). Prior work has made use of generative models for designing enhancer DNA sequences in specific cells [71]. Following Stark et al. [68], we report the perplexity over sequences as the main measure of performance. We also include results on Fréchet Biological Distance (FBD) with pre-trained classifiers provided in Dirichlet FM [68], cf.  §F.4.3. Nevertheless, those classifiers perform poorly on cell-type classification of Enhancer sequences, with test set accuracies of $11.5\%$ and $11.2\%$ on the Melanoma and FlyBrain datasets, respectively; thus, metrics derived from these are not representative of model quality, which we still included in §F.4.3 for transparency.
@@
Results. We report our results in Table 2, with FBD reported in Table 5. We observe that Fisher-Flow obtains significantly better performance than Dirichlet FM, which highlights its ability to fit the distribution of Melanoma and FlyBrain DNA enhancer sequences. Moreover, we also note that our method improves over the language model baseline on both datasets, which bolsters the belief that Fisher-Flow can be used in similar settings to those of autoregressive models.
4.4 De novo molecule generation
Table 3: Results on QM9. Higher is better. The baseline numbers are taken from the cited papers. The numbers reported for FlowMol are those for the uniform distribution and end-point parameterisation. Our numbers are for $1{,}000$ molecules.
Method Atoms S (%) Mols Val (%) Mols. S (%)
Fisher-Flow (ours) $98.6$ $95.3$ $88.2$
JODO [44] $99.4$ $98.9$ $98.7$
EquiFM [67] $99.4$ $94.4$ $93.2$
FlowMol [29] $98.9$ $96.9$ $84.2$
Figure 4: Generated molecules using Fisher-Flow on QM9.
In this experiment, we evaluate Fisher-Flow’s ability to generate molecules unconditionally, a.k.a. “de novo”. The difficulty in this task is that we are interested in generating the positions of the molecules, their atom types, their charges, and the bonds between these, resulting in a high dimensional space with both discrete and continuous data $({\mathbb{R}}^{d})^{n}\times(\Delta^{a})^{n}\times(\Delta^{c})^{n}\times(\Delta^{e})^{n^{2}}$, where $n\in\mathbb{N}^{\star}$ is the number of atoms, $a$ possible atom types, $c$ charges, and $e$ bonds. We train our model over the QM9 dataset [61, 60]. We report the percentage of stable atoms within molecules, valid molecules, and stable molecules. Our implementation is mostly based on that of [29].
Results. We report our results in Table 3. We also provide some qualitative examples in Figure 4. As we can see, Fisher-Flow compares well on all metrics to Simplex-Flow on all metrics. Nonetheless, it must be reported that the latter, trained with a Gaussian prior, endpoint parameterisation and cosine time schedule performed substantially better than both flow-based methods, closing the gap with the other baselines. It is likely that a more extensive exploration of priors, time parameterisations and other hyperparameters would increase Fisher-Flow’s performance.
4.5 Language modelling
Table 4: Test perplexities on the LM1B dataset. All baselines are taken from concurrent work MDLM by Sahoo et al. [62]. Best diffusion or flow-matching method is in bold font.
Method Parameters PPL ($\downarrow$)
Diffusion BERT-Mouth 110M $\leq 142.89$
D3PM (absorb) 70M $\leq 77.50$
Diffusion-LM 80M $\leq 118.62$
DiffusionBert 110M $\leq 63.78$
SEDD (33B tokens) 110M $\leq 32.79$
AR Transformer (33B tokens) 110M $22.32$
Transformer (327B tokens) 110M $20.86$
DM/FM MDLM (33B tokens) 110M $\leq 27.04$
Fisher-Flow (33B tokens) (ours) 110M $\mathbf{\leq 26.51}$
MDLM (327B tokens) 110M $\leq 23.00$
Fisher-Flow (327B tokens) (ours) 110M $\mathbf{\leq 22.42}$
Finally, we test the language modelling capabilities of Fisher-Flow. To do so, we train the model on the LM1B dataset [24], a large language modelling dataset containing about $800{,}000$ words. For this experiment, we extend Fisher-Flow to a masked path as is done by [62, 65]: we define the probability path as $p_{t}=\kappa_{t}p_{M}+(1-\kappa_{t})p_{\mathrm{unif}}$, where $\kappa:[0,1]\to[0,1]$ is a noise scheduler. Here, $p_{M}$ is the Fisher-Rao geodesic between the target, $x_{0}$, and the designated mask token $M$, while $p_{\mathrm{unif}}$ is also a Fisher-Rao geodesic between a sample from a uniform distribution and $x_{0}$. It is thus a convex combination of probability paths. Using a denoising architecture enables us to rewrite the original loss as a weighted negative log likelihood $-\mathbb{E}[\log p(x_{0}\mid x_{t})]$. This allows us to calculate an upper bound on the test perplexity, a natural evaluation metric for language modelling [62, 65].
Results. The results are given in Table 4. As one can observe, using the Fisher-Rao metric enables better performance than MDLM. Yet, the gap with auto-regressive methods is still significant.
@@ Acknowledgements
We thank Alexander Tong for his generous time, help, and guidance in helping with the language modeling experiments. OD is supported by both Project CETI and Intel. MP is supported by CenIA and by Chilean Fondecyt grant n. 1210426. AJB is partially supported by an NSERC Post-doc fellowship. This research is partially supported by EPSRC Turing AI World-Leading Research Fellowship No. EP/X040062/1 and EPSRC AI Hub on Mathematical Foundations of Intelligence: An "Erlangen Programme" for AI No. EP/Y028872/1
@@ Appendix B Geometry of the Simplex
    Then a restatement of Chentsov’s theorem [74, Thm. 11.1], [11, Thm. 1.2] is that if a sequence of Riemannian metrics $g_{d}$ over $\Delta^{d}$ defined for $d\geq 2$ satisfies the property that for any $1\leq d^{\prime}\leq d$ any Markov morphism $M:\Delta^{d^{\prime}}\to\Delta^{d}$ is an isometry with respect to metrics $g_{d^{\prime}},g_{d}$, then there exists $C>0$ such that each of the $g_{d}$ is $C$ times the Fisher-Rao metric on $\Delta^{d}$.
@@
For further details see Villani [76, Ch. 13].
@@
Step 1. Note that by a small modification of the proof, we can apply Villani [76, Thm. 10.42] to $\Delta^{d}$ with cost equal to $\mathbb{D}_{\rm{KL}}$, and obtain that the $W_{\rm{KL}}$-distance between an admissible competitor $p(\theta+\delta\theta)$ in Eq. 15 and $p(\theta_{n})$ is realised by a transport plan $T^{\delta\theta}$, such that we have $p(\theta+\delta\theta)=T^{\delta\theta}_{\#}p(\theta)$. By definition of $W_{\rm{KL}}$ and due to Chebyshev’s inequality, for all $C>0$, the set of points $S_{C}$ that $T^{\delta\theta}$ moves by more than $C\epsilon$ in $\mathbb{D}_{\rm{KL}}$-distance has $p(\theta)$-measure not larger than $1/C$. Furthermore, $T^{\delta\theta}$ is uniformly bounded over $\Delta^{d}_{c}\setminus S_{C}$ by our initial hypothesis. By approximating this transport plan by a flow (one can adapt the ideas from e.g., Santambrogio [63, Thm. 4.4] for this contstruction) over $S_{C}$, we can find a vector field $v^{\delta\theta}$ such that $v^{\delta\theta}(p_{\omega})=\frac{1}{\epsilon}\log_{p_{\omega}}(T^{\delta\theta}(p_{\omega}))+o_{\epsilon}(|\delta\theta|)$ for $p_{\omega}\in S_{C}$, with error uniformly bounded in $p_{\omega}\in{\mathcal{M}}$. We then extend $v^{\delta\theta}$ arbitrarily outside $S_{C}$. This procedure associates to each small enough change $\delta\theta$ a vector field $v_{\delta\theta}\in T_{p(\theta)}{\mathcal{P}}$ which whose time-$\epsilon$ flow, denoted $\phi_{v_{\delta\theta}}(t=\epsilon,\cdot)$ pushes measure $p(\theta)$ to a measure approaching $p(\theta+\delta\theta)$ in the limit $\epsilon\to 0,C\to\infty$.
@@ Appendix D Optimal Transport proofs
For point 1, we can use Villani [76, Thm. 10.28] (the simpler Villani [76, Thm. 10.41] also applies, with the minor modification that we work on a manifold with boundary). To verify its conditions, note that ${\mathcal{M}}^{d}\subset\mathbb{S}_{+}$ is a subset of a Riemannian manifold and has $(d-1)$-dimensional measure, and that cost $c(x,y)=d^{2}(x,y)$ is convex, thus it has unique superdifferential and $\nabla_{x}c(x,\cdot)$ is injective, as required.
For points 2 and 3, we note that by Villani [76, Cor. 7.22] (see also McCann [54]), in general Polish spaces displacement interpolants as given by Eq. 19 and Eq. 20, coincide with Wasserstein geodesics.
A simplified version of the proof of 4. is present in Santambrogio [63, Prop. 5.30]. For the general case, we can use Villani [76, Thm. 10.28], in particular eq. (10.20) therein. Note that for $c(x,y)=d^{2}(x,y)$, as indicated in Example 10.36 this equation corresponds to the equation of geodesics in the underlying manifold. Then we just note that $u_{t}$ is the velocity field of a constant speed geodesic.
Point 5 is a special case of Villani [76, Thm. 7.21], see also Granieri [33].
@@ E.1 Dirichlet Flow matching
where $\tilde{I}_{x}(a,b)=\frac{\partial}{\partial a}I_{x}(a,b)$ is the derivative of the regularised incomplete beta function [68, Appendix A.1] and $C\propto 1/t$ as in regular linear flow matching.